Concentration time data were collected for the following reaction at 30 oC
A + B --> 2C
Concentration time data at T = 30oC
Time (s) [A] (mol L-1)
0 0.65
15 0.52
30 0.35
45 0.28
60 0.22
75 0.14
90 0.095


(i) Calculate the average rate of consumption of A in the first 15 seconds of reaction.
(ii) Calculate the average rate of production of C in the first 15 seconds of reaction.
(iii) Calculate the average rate of consumption of A in the last 15 seconds of the reaction.
(iv) Explain the difference between the rates of consumption calculated in (i) and that in (iii)
Answer #1

Rate law can be written as follows.

Rate = -d[A]/dt = 1/2 d[C]/dt

Where -d[A]/dt is the rate of consumption of A.

And d[C]/dt is the rate of production of C.

(i) In the first 15 seconds of reaction, -d[A]/dt = (0.65-0.52) M/(15-0) s = 8.67*10-3 M/s

(ii) In the first 15 seconds of reaction, d[C]/dt = 2 * (8.67*10-3 M/s) = 1.73*10-2 M/s

(iii) In the last 15 seconds of reaction, -d[A]/dt = (0.14-0.095) M/(90-75) s = 3*10-3 M/s

(iv) The rate of sonsumption of A in the initial stage (first 15 seconds) is faster than that in the final stage (last 15 seconds).

i.e. 8.67*10-3 M/s > 3*10-3 M/s